OpenStudy (anonymous):

Hi All, I have a greatest common factor question. As usual, there is a question in the book that doesn't have not 1 example in the book:( How would I factor this polynomial using the gcf: 9x + 9? I know the answer is 8(x+1), that is, if the book is right, LOL!! I just need to know how to get there;) Thanks in advance, Denise:)

OpenStudy (anonymous):

oops, wrong answer, sorry, it's 8x + 1 8(x + 1) so sorry:(

OpenStudy (anonymous):

OMG, I've lost my marble, sec,

OpenStudy (anonymous):

8(x + 1) = 8*x + 8*1 = 8x + 8 8x + 8 != 9x + 9 Book is wrong or you mistyped, lol.

OpenStudy (anonymous):

8x + 1 = 8(x + 1)

OpenStudy (anonymous):

And just wanna say that I used != to mean doesn't always equal

OpenStudy (anonymous):

And that still doesn't work unless you're trying to solve for x and not factor. 8x + 1 = 8(x + 1) 8x + 1 = 8*x + 8*1 by the distributive property of multiplication over addition 8x + 1 = 8x + 8 Subtract 8x, 1 isn't 8 Not the same.

OpenStudy (anonymous):

I heard today that this 300 dollar book is going to have to be replaced because the instructors have found errors in it:( omg, my life is over, LOL

OpenStudy (anonymous):

Ok, how about this. How would I factor out(gcf) for 9x + 9?

OpenStudy (anonymous):

Let's just factor 9x + 9. What multiplies to 9x? 3, 9 and x, right? 9*x = 9x and 3*3*x = 9x What multiplies to 9? 1, 3, and 9. 1*9 = 9 and 3*3 = 9. What numbers do they have in common. Just 3 and 9. Which is the largest? 9. This means we can "take a 9 out" of both of those two things. Imagine I had a number 12. 12 is 2*2*3. If I "took a two out", I would get 2*3. "Taking a two out" of 12 gets 6. Taking a 9 out of 9x? 9*x. Take the 9 out, you're left with x. 9 out of 9? 1. Now we just put the number we took out (9) in front of the numbers that are left. 9(x + 1) Use the distributive property to check if you're right.

OpenStudy (anonymous):

Ok, I get it. I can go back up and use your equation sequence, that is how he has us do it:) Thank you soooooooooooooo much!!! Denise;)

OpenStudy (anonymous):

I don't see how the 8-8 on the right side equals 0 but the answer shows it as 1. I must have forgotten a rule like if the answer is 0 it always = 1??

OpenStudy (anonymous):

lol, i MEAN the 9-9 =0, LOL!! I'll be going back to first grade Spring term:(

OpenStudy (anonymous):

How does 9 out of 9 = 1, help:(

OpenStudy (radar):

It is more like 9 divided by 9 getting 1, or factor of 9 is also 9 and 1, as well as 3 * 3.